Your company buys from a supplier who offers credit terms of 3/10 net 130. Discuss whether your company should or should not pay cash for the goods it buys if it can borrow funds from the bank at 10% per annum. (Use a 360-day year for your computations.)
Answer #1

The credit terms "3 / 10 net 130" can be read as 3 % discount if paid within 10 days, else the total amount is due in 130 days

Formula for cost of credit is as below

Cost of credit = (1 + (discount %)/(1- discount %))^(360 / days past discount) - 1

Using the above parameters, cost of credit comes out to be

Cost of credit = (1 + 3%/97%)^(360/120)-1 = 9.57%

This cost of credit comes out to be less than the cost of funds from bank - Hence the company should go for trade credit instead of cash